The difference of two positive numbers is 8. the larger is Twice the smaller decreased by 7. find the two numbers
Answer by JulietG(1812)

x - y = 8
x = 2y - 7
Take the value of x from the second equation and substitute it into the first.
x - y = 8
(2y-7) - y = 8
y-7 = 8
y = 15
If y is 15, then x is 8 more than that, or 23.
23 = (2*15) - 7
23 = 30-7
23=23